If sec A+tanA=x , the secA=?

Grade:12

1 Answers

Chilukuri Sai Kartik
12 years ago

Dear Nitin,

Given that SecA+TanA=x------ Eq.1

We know that Sec2A-Tan2A=1

This is of the form a2-b2=(a+b)(a-b)

So (SecA-TanA)(SecA+TanA)=1

SecA-TanA=1/x   {From eq.1}

Lets consider this as Eq.2

 

Now adding Eq.1 and Eq.2, we get 2SecA=x+1/x

Hence SecA=(x+1/x)/2

 

So SecA=(x2+1)/2x

 

All the best!

 

Plz approve my answer if you are satisfied.